OpenStudy (anonymous):

iN JUST 12 HOURS ON FEBRUARY 21,1918 THE TEMERTURE GRANSVILLE, NORTH DAKOTA,ROSE FROM -33^F TO 50^F BY HOW MUCH DID THE TEMPERATURE CHANGE? I CAME UP WITH 50- -33=80 MY INSTRUCTOR SAID THIS WAS INCORRECT

OpenStudy (anonymous):

HE SAID I COMBINE THE TWO NUMBERS WRONG

Parth (parthkohli):

You just got one thing wrong.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 50 - (-33) = 50 + 33 = 83 \ne 80 }\)

Parth (parthkohli):

:)

OpenStudy (anonymous):

what does that sign before 80 mean ?

Parth (parthkohli):

I was just saying that 83 is not equal to 80 :)